# Dunkl approach to slice regular functions

### v3

@@ Abstract.
In this paper, we show that a Clifford algebra valued function is slice if and only if it is in the kernel of Dunkl-spherical Dirac operator and that a slice function $f$ is slice regular if and only if it is in the kernel of the Dunkl-Cauchy-Riemann operator of a certain parameter. Based on this observation and the inverse Dunkl intertwining operator, we present a new approach to construct a family of classical monogenic functions from a given holomorphic function.
@@ 1. Introduction
In the literature, there are mainly two approaches to generalize classical holomorphic function theory to higher dimensions. The first approach is the function theory of several complex variables (see e.g. [25]), while the second one revolves around the function theory of the Dirac operator, usually referred to as Clifford analysis [10]. Functions in the kernel of the Dirac operator are called monogenic. Over the last four decades, many results from complex analysis of one variable have been generalized within the latter framework. Unfortunately, the powers of the variables are not in the kernel of the classical Dirac operator. Since the 1990s, Leutwiler, Eriksson, and their collaborators have developed a modification of classical Clifford analysis that incorporates the powers of the variable into the kernel of the modified Dirac operator, see e.g. [26, 27]. Interestingly, their adjustment is closely related with hyperbolic geometry, see e.g. [17]. Around 2005, another generalization of classical Clifford analysis is introduced in [2] by employing the influential Dunkl operators [12], which has attracted a lot of attention in the last years. Some relations between these three kinds of monogenic functions have been discussed in e.g. [17, 24, 27]. See the right half part of Fig. 1 for a rough understanding.
@@
In 2006-2007, a new hyper-complex function theory, mainly studying so-called slice regular functions, was introduced by Gentili and Struppa in [21]. This theory has attracted considerable interest and experienced a rapid growth over the last years. In fact, the theory of slice regular functions of one variable nowadays is well-established and has found significant applications in the study of quaternionic quantum mechanics and spectral theory of several operators, see e.g. [6]. Importantly, convergent power series of the form $\sum_{n=0}^{\infty}x^{n}a_{n}$ are slice regular on an open ball centered at the origin. There is a huge amount of publications on this topic. Without claiming completeness, we refer to [5, 7, 21, 20].
@@
Fueter’s theorem (see e.g. [8, 15, 19, 30]) establishes a link between slice regular (or slice monogenic) functions and monogenic functions. These two theories are also intertwined through the Radon transform [4]. Besides those links, the hypermonogenic function theory, the Dunkl-Clifford analysis, the slice regular function theory were all developed independently in some sense.
@@
In this paper, we provide a link between these theories giving a new characterization of slice and slice regular functions using the Dunkl-spherical Dirac and the Dunkl-Cauchy-Riemann operators. More precisely, we find that a function defined in an axially symmetric open set is slice if and only if it is in the kernel of the Dunkl-spherical Dirac operator (Proposition 3.3) and it is slice regular if it is in the kernel of both the Dunkl-spherical Dirac operator and the Dunkl-Cauchy-Riemann operator (Theorem 3.6), with an appropriate condition on the multiplicity function. By this observation, we introduce a new approach in Section 5 to construct a family of classical monogenic functions from a given slice regular function based on the inverse Dunkl intertwining operator [14], whereas Fueter’s famous theorem only yields one axial monogenic function at one time. Moreover, our approach maps each $x^{n}$ homogeneously to monogenic polynomials of the same degree and doesn’t significantly differ in even and odd dimensions. The formula bears a resemblance to the well-known Cauchy-Kovalevskaya extension [10]. We would also like to remark that the structure of the inverse Dunkl intertwining operator is much simpler than the Dunkl intertwining operator, as the latter one is explicitly known only in some special cases, see e.g. [9, 14]. The connection with hyperbolic harmonic functions offers a local way to study slice regular functions, as shown in Theorem 3.8. Moreover, we find that the functions produced in the Fueter’s scheme are also in the kernel of certain Dunkl-Cauchy-Riemann operators. The Laplace operator in this scheme acts as a shift operator on the index $\gamma_{\kappa}$, see Section 4. These relations are summarized in Figure 1.
@@ 2.1. Dunkl operators
where $\langle\alpha,x\rangle:=\sum_{j=0}^{m}\alpha_{j}x_{j}$ is the standard Euclidean inner product and $|x|:=\sqrt{\langle x,x\rangle}$. A finite set $\mathscr{R}\subset\mathbb{R}^{m+1}\backslash\{0\}$ is called a root system if $\mathscr{R}\cap\mathbb{R}\cdot\alpha=\{\alpha,-\alpha\}$ and $\sigma_{\alpha}\mathscr{R}=\mathscr{R}$ for all $\alpha\in\mathscr{R}$. The dimension of ${\rm span}_{\mathbb{R}}\mathscr{R}$ is called the rank of $\mathscr{R}$. For a given root system $\mathscr{R}$, the reflections $\sigma_{\alpha}$, $\alpha\in\mathscr{R}$, generate a finite group $W\subset{\rm O}(m+1)$, called the Coxeter-group associated with $\mathscr{R}$. It is known that each root system can be expressed as a disjoint union, i.e. $\mathscr{R}=\mathscr{R}_{+}\cup(-\mathscr{R}_{+})$, where $\mathscr{R}_{+}$ and $-\mathscr{R}_{+}$ are separated by a hyperplane through the origin. We fix the positive subsystem $\mathscr{R}_{+}$. From now on, we assume that $\mathscr{R}$ is normalized, meaning that $\langle\alpha,\alpha\rangle=2$ for every $\alpha\in\mathscr{R}$. A function $\kappa:\mathscr{R}\rightarrow\mathbb{R}$ on a root system $\mathscr{R}$ is called a multiplicity function if it is invariant under the action of the associated reflection group $W$. For later use, we introduce the index $\gamma_{\kappa}:=\sum_{\alpha\in\mathscr{R}_{+}}\kappa(\alpha).$ The number $(m+1)+2\gamma_{\kappa}$ is the generalized homogeneous dimension, see [14].
@@ 2.2. Dunkl-Clifford analysis
$\begin{split}\left\{\begin{array}[]{ll}e_{j}\,e_{k}=-e_{k}\,e_{j},&1\leq j\neq k\leq m,\\ e_{j}^{2}=-1,&j=1,\ldots,m,\\ \end{array}\right.\end{split}$
where $(e_{1},e_{2},\ldots,e_{m})$ is the orthonormal basis of the Euclidean space $\mathbb{R}^{m}$. In the following, any vector $(x_{1},x_{2},\ldots,x_{m})\in\mathbb{R}^{m}$ will be identified with $\underline{x}=\sum_{j=1}^{m}x_{j}e_{j}$. The Clifford product of two vectors splits into a scalar part and a bivector as follows,
@@
while $D_{\kappa}$ reduces to the generalized Cauchy-Riemann operator ¯$\overline{\partial}_{x}:=\partial_{x_{0}}+\underline{D}.$
@@
(2.5) $\begin{split}\left\{\begin{array}[]{ll}\widetilde{\Gamma}_{\underline{\omega}}(M_{j})=-jM_{j},\\ \widetilde{\Gamma}_{\underline{\omega}}(\underline{x}M_{j})=(2\gamma_{\kappa}+j+m-1)\underline{x}M_{j}.\\ \end{array}\right.\end{split}$
@@
(2.6) $\begin{split}\left\{\begin{array}[]{ll}\widetilde{\Gamma}_{\underline{\omega}}f(r)=0,\\ \widetilde{\Gamma}_{\underline{\omega}}(\underline{\omega}f(r))=\widetilde{\Gamma}_{\underline{\omega}}(\underline{\omega})f(r),\\ \widetilde{\Gamma}_{\underline{\omega}}(\underline{\omega})=(2\gamma_{\kappa}+m-1)\underline{\omega}.\end{array}\right.\end{split}$
@@ 2.3. Slice regular functions
where $\overline{z}$ is the complex conjugation of $z$, and $\overline{(x\otimes z)}:=x\otimes\overline{z}$, for any $x\in\mathbb{R}_{0,m}$, $z\in\mathbb{C}$. Note that, if $F=\alpha+i\beta$, it is equivalent to requiring $\alpha(\overline{z})=\alpha(z)$ and $\beta(\overline{z})=-\beta(z)$. Let $\mathbb{S}^{m-1}$ be the $(m-1)$-dimensional unit sphere in $\mathbb{R}^{m}$. Then for any vector $\underline{\omega}\in\mathbb{S}^{m-1}$, it is easy to verify that $\underline{\omega}^{2}=-1$. Given $\underline{\omega}\in\mathbb{S}^{m-1}$, let $\mathbb{C}_{\underline{\omega}}={\rm span}\{1,\underline{\omega}\}$ be the complex slice of $\mathbb{R}^{m+1}$ generated by $\underline{\omega}$ and define the -algebra isomorphism $\phi_{\underline{\omega}}:\mathbb{C}\ni a+ib\mapsto a+\underline{\omega}b\in\mathbb{C}_{\underline{\omega}}$. We can define the circularization of a symmetric set $D$ as
@@
¯$\overline{\partial}_{\underline{\omega}}f(x_{0}+\underline{\omega}r):=\frac{1}{2}\left(\frac{\partial}{\partial x_{0}}+\underline{\omega}\frac{\partial}{\partial r}\right)f_{\underline{\omega}}(x_{0}+\underline{\omega}r)=0,$
@@
Furthermore, we introduce two notions related with slice functions, called the spherical value $f_{s}^{\circ}(x)$ and the spherical derivative $f_{s}^{\prime}(x)$ of $f$ in literature. They are defined as
(2.8) $\begin{split}\left\{\begin{array}[]{ll}f_{s}^{\circ}(x):=(f(x)+f(x^{c}))/2,\\ f_{s}^{\prime}(x):={\rm Im}(x)^{-1}(f(x)-f(x^{c}))/2,\\ \end{array}\right.\end{split}$
@@
defined on the class $\mathcal{C}^{1}(\Omega_{D}\backslash\mathbb{R})$ of $\mathbb{R}_{0,m}$-valued functions. The operator $\overline{\theta}$ is obtained by dividing the operator in [3] by $|\underline{x}|^{2}$. It was found that slice regularity can be characterized as follows:
@@ 3. Characterization of slice regular functions
where $\Psi$ is defined in (2.4). Thus we have
@@
    $\rho_{\underline{\alpha}}f(\sigma_{\underline{\alpha}}\underline{x})=\rho_{\underline{\alpha}}f(\underline{x})$;
@@
    $\rho_{\underline{\alpha}}f(\sigma_{\underline{\alpha}}\underline{x})=\dfrac{f(\sigma_{\underline{\alpha}}\underline{x})-f(\sigma_{\underline{\alpha}}^{2}\underline{x})}{\langle\underline{\alpha},\sigma_{\underline{\alpha}}\underline{x}\rangle}=\dfrac{f(\sigma_{\underline{\alpha}}\underline{x})-f(\underline{x})}{-\langle\underline{\alpha},\underline{x}\rangle}=\rho_{\underline{\alpha}}f(\underline{x}).$
@@
where $\mathcal{M}_{\ell}$ denotes the space of homogeneous polynomial of degree $\ell$ which are in the kernel of $\underline{D}_{\kappa}$. In other words, for any $P_{k}\in\mathcal{P}_{k}$ there exist unique $\mu_{\ell},\nu_{\ell}\in\mathcal{M}_{\ell}$ such that
@@
[…]
@@
where for any fixed $x_{0}$, $\mu^{(k)}_{\ell}(x_{0};\cdot),\nu^{(k)}_{\ell}(x_{0};\cdot)$ are homogeneous Dunkl monogenic polynomial of degree $\ell$ on $\mathbb{R}^{m}$. Furthermore, we remark that only the multiplicity function $\kappa\geq 0$ was considered in [28], however it can be generalized for all non-singular $\kappa$, as done in the scalar case [13].
Similar as the classical case, we have $\mu_{j}\perp\mu_{k}$ when $j\neq k$, and $\mu_{j}\perp\underline{x}\nu_{k}$ for any $j,k\in\mathbb{N}\cup\{0\}$ under the Fischer inner product. Now acting $\widetilde{\Gamma}_{\underline{\omega}}$ term by term in the decomposition (3.2), together with the eigenvalues of $\widetilde{\Gamma}_{\underline{\omega}}$ given in (2.5) and (2.6), we see that
$\begin{split}&\widetilde{\Gamma}_{\underline{\omega}}f(x_{0}+\underline{x})=\sum_{k=0}^{\infty}\sum_{j=0}^{[k/2]}|\underline{x}|^{2j}\left[\widetilde{\Gamma}_{\underline{\omega}}\left(\mu^{(k)}_{k-2j}(x_{0};\underline{x})\right)+\widetilde{\Gamma}_{\underline{\omega}}\left(\underline{x}\nu^{(k)}_{k-2j-1}(x_{0};\underline{x})\right)\right]\\ &=\sum_{k=0}^{\infty}\sum_{j=0}^{[k/2]}|\underline{x}|^{2j}\left[(2j-k)\mu^{(k)}_{k-2j}(x_{0};\underline{x})+(2\gamma_{\kappa}+k-2j-1+m-1)\underline{x}\nu^{(k)}_{k-2j-1}(x_{0};\underline{x})\right]\\ &=\sum_{k=0}^{\infty}\sum_{j=0}^{[k/2]}|\underline{x}|^{2j}\left[(2j-k)\mu^{(k)}_{k-2j}(x_{0};\underline{x})+(+k-2j-1)\underline{x}\nu^{(k)}_{k-2j-1}(x_{0};\underline{x})\right].\end{split}$
@@
$\widetilde{\Gamma}_{\underline{\omega}}f(x_{0}+\underline{x})=0\implies\left\{\begin{array}[]{ll}\mu^{(k)}_{k-2j}(x_{0};\underline{x})=0&\text{ if }k\text{ is odd or if }k\text{ is even and }j\neq k/2\\ \nu^{(k)}_{k-2j-1}(x_{0};\underline{x})=0&\text{ if }k\text{ is even or if }k\text{ is odd and }j\neq\frac{k-1}{2}\end{array}\right.$
@@
For the last claim, the second equality follows from Lemma 3.2 (3), while assuming $\widetilde{\Gamma}_{\underline{\omega}}f=0$, by Proposition 3.1 we have
@@
Let $\gamma_{\kappa}=(1-m)/2$. Then $\overline{\theta}$ and $D_{\kappa}$ agree on slice functions of class $\mathcal{C}^{1}(\Omega_{D}\backslash\mathbb{R})$.
@@
¯$\overline{\partial}_{x}f-\overline{\theta}f=2\gamma_{\kappa}f^{\prime}_{s}.$
@@
We can summarize Proposition 3.3 and Theorem 3.6 by
$\mathcal{S}(\Omega_{D})=\ker{\widetilde{\Gamma}_{\underline{\omega}}},\qquad\mathcal{S}\mathcal{R}(\Omega_{D})=\ker{\widetilde{\Gamma}_{\underline{\omega}}}\cap\ker{D_{\kappa}},$
where the second equality holds for slice domains.
@@
Suppose $f$ is slice regular, and denote $\underline{\omega}=\sum_{j=1}^{m}e_{j}\omega_{j}$ with $\omega_{j}\in\mathbb{R}$. Then $\alpha(x_{0},r)$ and $\omega_{j}\beta(x_{0},r)$ with $11\leq j\leq m-$ 1 are hyperbolic harmonic, i.e.
@@
while $\omega_{m}\beta(x_{0},r)$ is the eigenfunction of the hyperbolic Laplace operator corresponding to the eigenvalue $(1-m)$, i.e.
@@
By the splitting lemma for slice monogenic functions in [5], we only need to consider when $\alpha$ and $\beta$ are real-valued.
@@
[…]
@@
$\Delta_{\kappa}=\overline{D_{\kappa}}D_{\kappa}f=0.$
@@
Combining with (1), we get
(3.4) $\begin{split}\Delta_{\kappa}\,\underline{\omega}\,\beta(x_{0},r)=\left(\sum_{j=0}^{m-1}\partial_{x_{j}}^{2}+T_{m}^{2}\right)\,\underline{\omega}\,\beta(x_{0},r)=0.\end{split}$
@@
$\begin{split}T_{m}^{2}f=\left\{\begin{array}[]{ll}\displaystyle\partial_{x_{m}}^{2}+\frac{1-m}{x_{m}}\partial_{x_{m}}f,&\mbox{when $f$ is even on $x_{m}$},\\ \displaystyle\partial_{x_{m}}^{2}+\frac{1-m}{x_{m}}\partial_{x_{m}}f-\frac{1-m}{x_{m}^{2}}f,&\mbox{when $f$ is odd on $x_{m}$}.\\ \end{array}\right.\end{split}$
The above (3.4) together with the fact that $x_{j}\beta(x_{0},r)$ is even on $x_{m}$ when $11\leq j\leq m-$ 1 and odd when $j=m$, yields the claimed relations of $\beta$, which concludes the proof. ∎

### v4

@@ Abstract.
In this paper, we establish a connection between Dunkl analysis and slice analysis in the setting of Clifford algebras. Specifically, we show that a Clifford algebra-valued function is slice if, and only if, it belongs to the kernel of the Dunkl-spherical Dirac operator and that a slice function is slice regular if, and only if, it lies in the kernel of the Dunkl-Cauchy-Riemann operator for a suitable parameter. Based on this correspondence and the inverse Dunkl intertwining operator, we propose a new method to construct a family of classical monogenic functions from a given holomorphic function, in the spirit of Fueter’s theorem.
@@ 1. Introduction
In the literature, there are two main approaches to generalize classical holomorphic function theory to higher dimensions. The first is the function theory of several complex variables (see e.g. [25]), while the second revolves around the function theory of the Dirac operator, usually referred to as Clifford analysis [10]. Functions in the kernel of the Dirac operator are called monogenic. Over the past four decades, many results from complex analysis of one variable have been generalized within the latter framework. Unfortunately, the powers of the variables are not in the kernel of the classical Dirac operator. Since the 1990s, Leutwiler, Eriksson, and their collaborators have developed a modification of classical Clifford analysis that incorporates the powers of the variable into the kernel of the modified Dirac operator, see e.g. [26, 27]. Interestingly, their adjustment is closely related to hyperbolic geometry, see e.g. [17]. Around 2005, another generalization of classical Clifford analysis was introduced in [2] by employing the influential Dunkl operators [12], which has attracted significant attention in recent years. Some relations between these three types of monogenic functions have been discussed in e.g. [17, 24, 27]. See the right half of Fig. 1 for a rough understanding.
@@
In 2006-2007, a new hypercomplex function theory, mainly studying so-called slice regular functions, was introduced by Gentili and Struppa in [21]. This theory has attracted considerable interest and experienced a rapid growth over the last years. In fact, the theory of slice regular functions of one variable nowadays is well-established and has found significant applications in quaternionic quantum mechanics and spectral theory of several operators, see e.g. [6]. Importantly, a convergent power series of the form $\sum_{n=0}^{\infty}x^{n}a_{n}$ defines a slice regular function on an open ball centered at the origin. There is a vast amount of publications on this topic. Without claiming completeness, we refer to [5, 7, 21, 20].
@@
Fueter’s theorem (see e.g. [8, 15, 19, 30]) establishes a link between slice regular (or slice monogenic) functions and monogenic functions. These two theories are also connected through the Radon transform [4]. Besides those links, the theories of hypermonogenic functions, Dunkl-Clifford analysis and slice regular functions have all been developed independently to some extent.
@@
In this paper, we establish a link between these theories providing a new characterization of slice and slice regular functions using the Dunkl-spherical Dirac and the Dunkl-Cauchy-Riemann operators. More precisely, we show that a function defined in an axially symmetric open set is slice if and only if it lies in the kernel of the Dunkl-spherical Dirac operator (Proposition 3.3) and, for slice domains, it is slice regular if it is in the kernel of both the Dunkl-spherical Dirac operator and the Dunkl-Cauchy-Riemann operator (Theorem 3.6), under an appropriate condition on the multiplicity function. By this observation, we introduce a new approach in Section 5 to construct a family of classical monogenic functions from a given slice regular function, based on the inverse Dunkl intertwining operator [14]. In contrast, Fueter’s famous theorem only produces a single axial monogenic function at a time. Moreover, our approach maps each $x^{n}$ homogeneously to monogenic polynomials of the same degree and does not exhibit significant differences between even and odd dimensions. The resulting formula bears a resemblance to the well-known Cauchy-Kovalevskaya extension [10]. We also remark that the structure of the inverse Dunkl intertwining operator is much simpler than the Dunkl intertwining operator, as the latter one is explicitly known only in some special cases, see e.g. [9, 14]. The connection with hyperbolic harmonic functions provides a local approach to study slice regular functions, as shown in Theorem 3.8. Moreover, we find that the functions produced via Fueter’s scheme also lie in the kernel of certain Dunkl-Cauchy-Riemann operators. The Laplace operator in this scheme acts as a shift operator on the index $\gamma_{\kappa}$, see Section 4. These relations are summarized in Figure 1.
@@ 2.1. Dunkl operators
[…]
@@ 2.2. Dunkl-Clifford analysis
[…]
where $(e_{1},e_{2},\ldots,e_{m})$ is an orthonormal basis of the Euclidean space $\mathbb{R}^{m}$. In the following, any vector $(x_{1},x_{2},\ldots,x_{m})\in\mathbb{R}^{m}$ will be identified with $\underline{x}=\sum_{j=1}^{m}x_{j}e_{j}$. The Clifford product of two vectors splits into a scalar part and a bivector as follows,
@@
while $D_{\kappa}$ reduces to the generalized Cauchy-Riemann operator $\overline{\partial}_{x}:=\partial_{x_{0}}+\underline{D}.$
@@
[…]
@@
[…]
@@ 2.3. Slice regular functions
where $\overline{z}$ is the complex conjugation of $z$, and $\overline{(x\otimes z)}:=x\otimes\overline{z}$, for any $x\in\mathbb{R}_{0,m}$, $z\in\mathbb{C}$. Note that, if $F=\alpha+i\beta$, it is equivalent to requiring $\alpha(\overline{z})=\alpha(z)$ and $\beta(\overline{z})=-\beta(z)$. Let $\mathbb{S}^{m-1}$ be the (m1)(m-1)-dimensional unit sphere in $\mathbb{R}^{m}$. Then for any vector $\underline{\omega}\in\mathbb{S}^{m-1}$, it is easy to verify that $\underline{\omega}^{2}=-1$. Given $\underline{\omega}\in\mathbb{S}^{m-1}$, let $\mathbb{C}_{\underline{\omega}}={\rm span}\{1,\underline{\omega}\}$ be the complex slice of $\mathbb{R}^{m+1}$ generated by $\underline{\omega}$ and define the -algebra isomorphism $\phi_{\underline{\omega}}:\mathbb{C}\ni a+ib\mapsto a+\underline{\omega}b\in\mathbb{C}_{\underline{\omega}}$. We can define the circularization of a symmetric set $D$ as
@@
$\overline{\partial}_{\underline{\omega}}f(x_{0}+\underline{\omega}r):=\frac{1}{2}\left(\frac{\partial}{\partial x_{0}}+\underline{\omega}\frac{\partial}{\partial r}\right)f_{\underline{\omega}}(x_{0}+\underline{\omega}r)=0,$
@@
Furthermore, we introduce two notions related with slice functions, called the spherical value $f_{s}^{\circ}(x)$ and the spherical derivative $f_{s}^{\prime}(x)$ of $f$. They are defined as
[…]
@@
defined on the class $\)\mathcal{C}^{1}(\Omega_{D}\backslash\mathbb{R}$) of $\mathbb{R}_{0,m}$-valued functions. The operator $\overline{\theta}$ is obtained by dividing the operator in [3] by $|\underline{x}|^{2}$. It was found that slice regularity can be characterized as follows:
@@ 3. Characterization of slice regular functions
where $\Psi$ is defined in (2.4). Thus, we have
@@
    $\rho_{\underline{\alpha}}(f\circ\sigma_{\underline{\alpha}})=-\rho_{\underline{\alpha}}f$;
@@
    Since $\sigma_{\underline{\alpha}}^{2}=\operatorname{Id}$, for any $\underline{x}$ we have
    $\rho_{\underline{\alpha}}(f\circ\sigma_{\underline{\alpha}})(\underline{x})=\dfrac{f(\sigma_{\underline{\alpha}}\underline{x})-f(\sigma_{\underline{\alpha}}^{2}\underline{x})}{\langle\underline{\alpha},\underline{x}\rangle}=\dfrac{f(\sigma_{\underline{\alpha}}\underline{x})-f(\underline{x})}{\langle\underline{\alpha},\underline{x}\rangle}=-\rho_{\underline{\alpha}}f(\underline{x}).$
@@
where $\mathcal{M}_{\ell}$ denotes the space of homogeneous polynomials of degree $\ell$ which are in the kernel of $\underline{D}_{\kappa}$. In other words, for any $P_{k}\in\mathcal{P}_{k}$ there exist unique $\mu_{\ell},\nu_{\ell}\in\mathcal{M}_{\ell}$ such that
@@
$\begin{split}\widetilde{\Gamma}_{\underline{\omega}}f(x_{0}+\underline{x})=&\,\widetilde{\Gamma}_{\underline{\omega}}\alpha(x_{0},r)+\widetilde{\Gamma}_{\underline{\omega}}(\underline{\omega}\,\beta(x_{0},r))\\ =&\,(2\gamma_{\kappa}+m-1)\underline{\omega}\,\beta(x_{0},r)\\ =&\,0.\end{split}$
@@
where for any fixed $x_{0}$, $\mu^{(k)}_{\ell}(x_{0};\cdot),\nu^{(k)}_{\ell}(x_{0};\cdot)$ are homogeneous Dunkl monogenic polynomial of degree $\ell$ on $\mathbb{R}^{m}$. Furthermore, we remark that only the multiplicity function $\kappa\geq 0$ was considered in [28], however it can be generalized for all non-singular $\kappa$, as done in the scalar case [13]. Similar to the classical case, we have $\mu_{j}\perp\mu_{k}$ when $j\neq k$, and $\mu_{j}\perp\underline{x}\nu_{k}$ for any $j,k\in\mathbb{N}\cup\{0\}$ under the Fischer inner product. Now acting $\widetilde{\Gamma}_{\underline{\omega}}$ term by term in the decomposition (3.2), together with the eigenvalues of $\widetilde{\Gamma}_{\underline{\omega}}$ given in (2.5) and (2.6), we see that
$\begin{split}&\widetilde{\Gamma}_{\underline{\omega}}f(x_{0}+\underline{x})=\sum_{k=0}^{\infty}\sum_{j=0}^{[k/2]}|\underline{x}|^{2j}\left[\widetilde{\Gamma}_{\underline{\omega}}\left(\mu^{(k)}_{k-2j}(x_{0};\underline{x})\right)+\widetilde{\Gamma}_{\underline{\omega}}\left(\underline{x}\nu^{(k)}_{k-2j-1}(x_{0};\underline{x})\right)\right]\\ &=\sum_{k=0}^{\infty}\sum_{j=0}^{[k/2]}|\underline{x}|^{2j}\left[(2j-k)\mu^{(k)}_{k-2j}(x_{0};\underline{x})+(2\gamma_{\kappa}+k-2j-1+m-1)\underline{x}\nu^{(k)}_{k-2j-1}(x_{0};\underline{x})\right]\\ &=\sum_{k=0}^{\infty}\sum_{j=0}^{[k/2]}|\underline{x}|^{2j}\left[(2j-k)\mu^{(k)}_{k-2j}(x_{0};\underline{x})+(k-2j-1)\underline{x}\nu^{(k)}_{k-2j-1}(x_{0};\underline{x})\right].\end{split}$
@@
[…]
@@
For the last claim, the second equality follows from Lemma 3.2 (3), while assuming $\widetilde{\Gamma}_{\underline{\omega}}f=0$ and by Proposition 3.1 we have
@@
Let $\gamma_{\kappa}=(1-m)/2$. Then $\overline{\theta}$ and $D_{\kappa}$ agree on slice functions of class $\)\mathcal{C}^{1}(\Omega_{D}\backslash\mathbb{R}$).
@@
$\overline{\partial}_{x}f-\overline{\theta}f=2\gamma_{\kappa}f^{\prime}_{s}.$
@@
Suppose $f$ is slice regular, and denote $\underline{\omega}=\sum_{j=1}^{m}e_{j}\omega_{j}$ with $\omega_{j}\in\mathbb{R}$. Then $\alpha(x_{0},r)$ and $\omega_{j}\beta(x_{0},r)$ with $1\leq j\leq m-1$ are hyperbolic harmonic, i.e.
@@
while $\omega_{m}\beta(x_{0},r)$ is the eigenfunction of the hyperbolic Laplace operator corresponding to the eigenvalue (1m)(1-m), i.e.
@@
$\begin{split}&\left(x_{m}^{2}\Delta_{\mathbb{R}^{m+1}}-(m-1)x_{m}\partial_{x_{m}}\right)\alpha(x_{0},r)\\ =&\,\left(x_{m}^{2}\left(\partial_{x_{0}}^{2}+\partial_{r}^{2}+\frac{m-1}{r}\partial_{r}\right)-x_{m}^{2}\frac{(m-1)}{r}\partial_{r}\right)\alpha(x_{0},r)\\ =&\,x_{m}^{2}\left(\partial^{2}_{x_{0}}+\partial_{r}^{2}\right)\alpha(x_{0},r)\\ =&\,0.\end{split}$
@@
$\Delta_{\kappa}f=\overline{D_{\kappa}}D_{\kappa}f=0.$
@@
[…]
Now since $\alpha(x_{0},r)$ is even on $x_{m}$, the result in (1) is equivalent with
$\Delta_{\kappa}\,\alpha(x_{0},r)=\left(\sum_{j=0}^{m-1}\partial_{x_{j}}^{2}+T_{m}^{2}\right)\,\alpha(x_{0},r)=0.$
Together with $\Delta_{\kappa}f=0$, we get
(3.4) $\begin{split}&\Delta_{\kappa}\,\underline{\omega}\,\beta(x_{0},r)=\left(\sum_{j=0}^{m-1}\partial_{x_{j}}^{2}+T_{m}^{2}\right)\,(f-\alpha(x_{0},r))=0.\end{split}$
The equation (3.4) together with the fact that $x_{j}\beta(x_{0},r)$ is even on $x_{m}$ when $1\leq j\leq m-1$ and odd when $j=m$, yields the claimed relations for $\beta$, which concludes the proof. ∎